OpenStudy (anonymous):

solve by using the substitution method 5x-6y=16 7x+89=y

OpenStudy (anonymous):

5x - 6 (7x + 89) = 16

OpenStudy (anonymous):

Since you know that y = 7x + 89, you can substitute 7x + 89 for y in the first equation. Now you are only dealing with one variable rather than two.

OpenStudy (anonymous):

Once you solve for x, you can substitute that value into the second equation to determine the value of y.

OpenStudy (anonymous):

IM SO CONFUSED CAN U HELP ME SOLVE THIS IN A WAY I CAN LEARN IT PLEASE

OpenStudy (anonymous):

You are given two equations. (1) 5x-6y=16 (2) 7x+89=y Let's call the first one equation (1) and the second one equation (2) Note: equation (2) can also be written: y = 7x + 89

OpenStudy (anonymous):

Each equations has two variables. We do not want to work with equations that have two variables. So we will use substitution to get rid of one of the variables. Equation (2) tells us that y is equal to 7x + 89 So lets rewrite equation one, but instead of writing the y we will substitute 7x + 89. So here is the rewritten equation (1): 5x - 6 (7x + 89) = 16 Note: I wrote equation (1) but instead of the y, I plugged in 7x + 89.

OpenStudy (anonymous):

Now I will use the rewritten equation (1) to solve for x: 5x - 42x - 534 = 16 [I distributed the -6] -37x - 534 = 16 [I combined like terms] -37x = 550 [I added 534 to both sides of the equation] x = 550/-37 or -550/37 [This problem is especially tough because we end up with a large, negative fraction] Now that we know what x equals, we plug in that value in equation (2)

OpenStudy (anonymous):

y = 7x + 89 y = 7(-550/37) + 89 y = -3850/37 + 89 Plug that into your calculator to determine what y equals. This problem is quite messy. Are you sure you copied it correctly?